IR-UV RESONANT TWO-PHOTON IONIZATION SPECTROSCOPY OF Sc_{n}(BENZENE)_{m} (n =1, 2; m =1-3) AND Sc(BIPHENYL)

\underline{YUXIU LEI}, JUNG SUP LEE and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

The complexes of Sc_{n}(C_{6}H_{6})_{m} (n =1, 2; m =1-3, n≤ m) and Sc(C_{12}H_{10}) were synthesized in a laser-vaporization metal cluster beam source. The C-H stretching frequencies of these complexes were measured with IR-UV resonant two-photon ionization spectroscopy. The experimental C-H stretching frequencies were compared with those of the free ligands and with the B3LYP/6-311+G(d,p) calculations. From these comparisons, the ground electronic states were determined to be ^{4}A_{1} (C_{6v}) for the half-sandwich Sc(C_{6}H_{6}), ^{2}B_{3g}(D_{2h}) for the sandwich Sc(C_{6}H_{6})_{2} , ^{1}A(D_{2}) for the double sandwich Sc_{2}(C_{6}H_{6})_{3} and^{2}B_{1}(C_{2v}) for the clamshell Sc(C_{12}H_{10}). The conclusions from this work are in consistent with the previous ZEKE spectroscopic studies.^{a}^{,}^{b}

aB. R. Sohnlein, S. Li, and D.-S. Yang, J. Chem. Phys. 123 (21), 214306 (2005).

bB. R. Sohnlein, J. F. Fuller, and D.-S. Yang, J. Am. Chem. Soc. 128 (33), 10692 (2006).